OpenStudy (anonymous):

Solve using substitution. (2+(3/t))^2-(2+(3/t))=12

OpenStudy (anonymous):

\[(2+\frac{ 3 }{ t })^{2}-(2+\frac{ 3 }{ t })=12\]

OpenStudy (anonymous):

put \(x=2+\frac{3}{t}\)and solve \[x^2-x=12\] or \[x^2-x-12=0\] by factoring

OpenStudy (anonymous):

Awesome, thanks.